### v1

@@ 3.3 Applications of XAI for Operations in the Industry
The process industry is a subset of businesses that manufacture items from raw materials (not components) using formulae or recipes. Given the magnitude and dynamic nature of operations in the process sector, it becomes evident that the next great step ahead will be the capacity for people and AI systems to collaborate to ensure production stability and dependability [49]. AI systems must successfully inform the individuals who share the ecosystem about their objectives, intentions, and findings as the first step toward collaboration. In the future, people will work ”with” automation rather than ”around” it, thanks in part to the systematic approach to XAI.
This research [50] focuses on Explainable Artificial Intelligence (XAI) applications in the process industry. The research argues that current AI models are not transparent enough for process industry applications, and highlights the need for XAI models that can be understood by human experts. The main contribution is outlining the challenges and research needs for XAI in the process industry. The outcome is to develop XAI models that are safe, reliable, and meet the needs of human users in the process industry.
@@ 4 Future of Trustworthy (XAI)
The precise position of each XAI domain and how they relate to the human user are shown in Figure 13. The majority of AI system explanations that are given are usually static and only contain one message [66]. Understanding cannot be attained by explanations alone [67]. Because most existing XAI libraries lack user involvement and customization of explanations, users should be able to explore the system using interactive explanations to gain a better understanding of it. This is a promising research direction for extending the XAI field [67] and [66]. To improve human-machine cooperation and move beyond static explanations, a number of efforts have also been proposed.
Explainable Artificial Intelligence (XAI) has great promise for redefining the relationship between humans and AI systems as it stands at the nexus of technological innovation and societal integration. As AI technologies advance, it is more important than ever to ensure accountability and transparency. Within this framework, XAI becomes a crucial facilitator, entrusted with shedding light on the murky inner workings of AI models and cultivating user confidence. A wide range of breakthroughs are anticipated in XAI, from heightened model transparency and human-centric design principles to regulatory compliance requirements and the rise of hybrid AI systems. XAI approaches will place a high value on user-centric design, providing explanations that are both actionable and understandable. This will foster acceptance and confidence in AI systems.
Moreover, it is anticipated that regulatory frameworks would require the incorporation of XAI in essential applications, guaranteeing compliance with accountability and transparency norms. Future XAI systems will be distinguished by their contextual sensitivity and interactive explanations, which will enable users to interact with AI decisions in real time and adjust to a variety of situations. To guarantee that AI systems follow ethical standards and social values, as well as to democratize access to XAI technologies, efforts must be made to enhance digital literacy and address ethical challenges. Fundamentally, XAI’s success is predicated on its capacity to bridge the communication gap between AI systems and human users, fostering mutual respect, trust, and collaboration in an increasingly AI-dependent world.
@@
This study [68] offers a thorough analysis of Explainable Artificial Intelligence (XAI), focusing on two primary areas of inquiry: general XAI difficulties and research directions, as well as ML life cycle phases-based challenges and research directions. In order to shed light on the significance of formalism, customization of explanations, encouraging reliable AI, interdisciplinary partnerships, interpretability-performance trade-offs, and other topics, the study synthesizes important points from the body of existing literature. The primary contribution is the methodical synthesis and analysis of the body of literature to identify important problems and future directions for XAI research [68]. The research offers a thorough review of the current state of XAI research and provides insightful information for future studies and breakthroughs in the area by structuring the debate around general issues and ML life cycle phases. The primary finding of the study is the identification and clarification of 39 important points that cover a range of issues and potential avenues for future XAI research. The importance of conveying data quality, utilizing human expertise in model development, applying rule extraction for interpretability, addressing security concerns, investigating XAI for reinforcement learning and safety, and taking into account the implications of privacy rights in explanation are just a few of the many topics covered by these points. Furthermore, the paper indicates directions for further research and application by highlighting the potential contributions that XAI may make to a number of fields, including digital forensics, IoT, and 5G.
@@ 5 Conclusions
Explainable Artificial Intelligence (XAI) is gaining popularity in a range of fields due to its critical role in addressing critical issues connected to AI adoption. As AI systems become more integrated into society, transparency and interpretability become increasingly important. By offering tools to clarify how AI models make decisions, XAI helps users develop a sense of confidence and comprehension. XAI’s primary objective is to make AI models clear and intelligible. With the help of XAI, the general public will be able to peer inside the black box and comprehend the aspects that affect the AI’s decision-making process. The paper discusses the essential details of XAI and offers a comprehensive overview for a solid understanding. Furthermore, this article discusses in detail the three main application fields of XAI. Lastly, the authors attempt to outline the difficulties in applying XAI and suggest potential future paths.

### v2

@@ 3.3 Applications of XAI for Operations in the Industry
The process industry is a subset of businesses that manufacture items from raw materials (not components) using formulae or recipes. Given the magnitude and dynamic nature of operations in the process sector, it becomes evident that the next great step ahead will be the capacity for people and AI systems to collaborate to ensure production stability and dependability [50]. AI systems must successfully inform the individuals who share the ecosystem about their objectives, intentions, and findings as the first step toward collaboration. We can hope people to work ”with” automation rather than ”around” it, thanks in part to the systematic approach to XAI.
This research [51] focuses on XAI applications in the process industry. The research argues that current AI models are not transparent enough for process industry applications and highlights the need for XAI models that can be understood by human experts. The main contribution is outlining the challenges and research needs for XAI in the process industry. The outcome is to develop XAI models that are safe, reliable, and meet the needs of human users in the process industry.
@@ 4 Future of Trustworthy (XAI)
Figure 13 illustrates the precise location of each XAI domain and its relationship with the human user. According to [67], many explanations of AI systems tend to be static and convey only a single message. However, explanations alone do not facilitate true understanding [68]. To enhance comprehension, users should have the ability to explore the system through interactive explanations, as most existing XAI libraries currently lack options for user engagement and explanation customization. This represents a promising avenue for advancing the field of XAI [68] and [67]. Additionally, various efforts have been made to improve human-machine collaboration by moving beyond static explanations.
Explainable AI offers a way to improve how people interact with AI systems. As AI technology grows, it is crucial to ensure that these systems are accountable and transparent. XAI helps by clarifying how AI models work and building trust among users.
We can expect many new developments in XAI. These include making AI models more open, focusing on human-centered designs, ensuring compliance with regulations, and creating hybrid AI systems. XAI will prioritize designs that are easy for users to understand and provide clear explanations. This clarity will help build trust and encourage more people to use AI systems.
Regulatory frameworks are likely to require the use of XAI in important areas to ensure accountability and transparency. Future XAI systems will need to be sensitive to context and provide interactive explanations. This will allow people to engage with AI decisions in real time and adapt to different situations. We must also work to improve digital literacy and tackle ethical issues to ensure that AI systems follow moral principles and society’s values, making XAI technologies accessible to everyone. The success of XAI depends on its ability to bridge the communication gap between AI systems and human users, which encourages cooperation, mutual respect, and trust in an increasingly AI-driven world.
@@
This study [69] offers a thorough analysis of XAI, focusing on two primary areas of inquiry: general XAI difficulties and research directions, as well as ML life cycle phase-based challenges and research directions. In order to shed light on the significance of formalism, customization of explanations, encouraging reliable AI, interdisciplinary partnerships, interpretability-performance trade-offs, and other topics, the study synthesizes important points from the body of existing literature. The primary contribution is the methodical synthesis and analysis of the body of literature to identify important problems and future directions for XAI research [69]. The research offers a thorough review of the current state of XAI research and provides insightful information for future studies and breakthroughs in the area by structuring the debate around general issues and ML life cycle phases. The primary finding of the study is the identification and clarification of 39 important points that cover a range of issues and potential avenues for future XAI research. The importance of conveying data quality, utilizing human expertise in model development, applying rule extraction for interpretability, addressing security concerns, investigating XAI for reinforcement learning and safety, and taking into account the implications of privacy rights in explanation are just a few of the many topics covered by these points. Furthermore, the paper indicates directions for further research and application by highlighting the potential contributions that XAI may make to a number of fields, including digital forensics, IoT, and 5G.
@@ 5 Conclusions
XAI, or Explainable Artificial Intelligence, is becoming important in many industries because it helps solve key challenges with using AI. As AI becomes more common in our daily lives, understanding how it works is essential. XAI provides tools that help people see and understand how AI models make decisions. The main goal of XAI is to make these models easier to understand. It allows people to look inside the ”black box” of AI and see what affects its decisions. The paper gives a clear overview of the key parts of XAI. It also discusses three main areas where XAI can be applied. Finally, the authors talk about the challenges of using XAI and suggest possible future directions.